OpenStudy (selena2345):

You begin a hike in Death Valley,California at an elevation at an elevation of -86 meters. You hike to a point of elevation at 45 meters. What is your change in elevation ?

OpenStudy (anonymous):

It's xfinal - xinitial, so in this case, 45 - (-86) = 45+86 = 131m

OpenStudy (anonymous):

-86 + x = 45 x = 45 + 86 x = 131